On January 1, 2021, Packard Corporation leased equipment to Hewlitt Company. The lease term is 8 years. The first payment of $459,000 was made on January 1, 2021. Remaining payments are made on December 31 each year, beginning with December 31, 2021. The equipment cost Packard Corporation $2,741,434. The present value of the lease payments is $2,769,125. The lease is appropriately classified as a sales-type lease. Assuming the interest rate for this lease is 9%, what will be the balance reported as a liability by Hewlitt in the December 31, 2022, balance sheet? (Round your final answer to nearest dollar.)

Multiple Choice

  • $2,059,036.

  • $1,785,349.

  • $2,310,125.

  • $1,535,349.
Answer #1
Lease liability at inception 2769125
Less: Payment on January 1, 2021 459000
Lease liability at January 1, 2021 2310125
Interest for December 31,2021 207911 =2310125*9%
Lease liability reduction for December 31,2021 251089 =459000-207911
Lease liability at December 31,2021 2059036 =2310125-251089
Interest for December 31,2022 185313 =2059036*9%
Lease liability reduction for December 31,2022 273687 =459000-185313
Lease liability at December 31,2022 1785349 =2059036-273687
Option B $1,785,349 is correct